luvvdani1516:

Can someone please explain how I find the extraneous solution? I'm stuck and don't know what to do. I tried looking it up, but I don't understand what to do... SS below in comments

luvvdani1516:

1 attachment
surjithayer:

you are correct. x=-7 is an extraneous solution. |x-5|=2(x+1) when x=-7 |-7-5|=2(-7+1) |-12|=2(-6) 12=-12 which is not true ,hence x=-7 is an extraneous solution. solving x-5=-2(x+1) x-5=-2x-2 x+2x=-2+5 3x=3 x=3/3=1 when x=1 |x-5|=2(x+1) |1-5|=2(1+1) |-4|=2(2) 4=4 which is true. so x=1 is a solution.

luvvdani1516:

@surjithayer wrote:
you are correct. x=-7 is an extraneous solution. |x-5|=2(x+1) when x=-7 |-7-5|=2(-7+1) |-12|=2(-6) 12=-12 which is not true ,hence x=-7 is an extraneous solution. solving x-5=-2(x+1) x-5=-2x-2 x+2x=-2+5 3x=3 x=3/3=1 when x=1 |x-5|=2(x+1) |1-5|=2(1+1) |-4|=2(2) 4=4 which is true. so x=1 is a solution.
Ohhhh thank you!

surjithayer:

yw